OpenStudy (anonymous):

find the derivative of y with respect to x. y=tan^-1 (sqrt 3x)

OpenStudy (helder_edwin):

use the chain rule. you have 3 functions that are composed \[ f(x)=\arctan x=tan^{-1}x \] \[ g(x)=\sqrt{x} \] \[ h(x)=3x \] so \[ y=(f\circ g\circ h)(x) \] and the chain rule gives you \[\LARGE y'=(f'\circ g\circ h)(x)\cdot(g'\circ h)(x)\cdot h'(x) \] try it!

OpenStudy (anonymous):

or you can draw a triangle...

OpenStudy (anonymous):